You are studying tail length in squirrels. Short tail (S) is dominant to long tail (1). If you cross a homozygous dominant (S i need help solving this question
Answer #1

Answer-

According to the given question-

Here we have characteristics of tail length, where the short tail is dominant over the tall tail.

The allele for Short tail = S

The allele for long-tail = l

When we make a cross between the one parent with a homozygous short tail and another parent with homozygous tall, then in the F1 generation we get following offspring-

The genotype of homozygous short tail = SS

The genotype of homozygous long tail = ll

Cross between SS \times ll

l l
S Sl (Short tail ) Sl (Short tail )
S Sl (Short tail ) Sl(Short tail )

Then in the F1 generation, all the offspring is having short tail.

But the observed frequency of offspring n the F1 generation has given -

Short tail = 70

Long tail = 45

Total offspring = 70 + 45 = 115

Expested number of offspring in the F1 generation will be-

Short tail = 115

Long tail = 00

Chi square characteristics-

Phenotype Observed frequency (O) Expected frequency (E) O -E (O -E)2 (O -E)2 \div E
Short tail 70 115 70 - 115 = -45 2025 2025 \div 115 = 17.61
Long tail 45 00 45-0= 45 2025 2025 \div 0 = 0
Total = 115 Total = 115 X2 = 17.61

Here the calculated value of Chi square = X2 = 17.61

Degree of freedom = 2- 1= 1

The critical value for 1 degree of freedom at the ptobability of 0.05 = 3.841

Here the calculated value of Chi-square is more than the critical value of chi-square for 1 degree of freedom at the probability of 0.05 i.e. X2 Calculated > X2 Critical or 17. 61 > 3.841

Thus we reject the null hypothesis and it means that the observed trait is not according to the Mendelian law, and the observed value is different from the expected value.

Answer is (B)